if 24 oranges cost $3.50, how many oranges can you buy for $5.25

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)


5.25/(3.50/24)=36

Answer by stanbon(75887)

if 24 oranges cost $3.50, how many oranges can you buy for $5.25
------
Use a proportion::
#/cost = #/cost
=========================
x/525 = 24/350
----
x = 525(24/350)
---
x = 21(24/14)
----
x = 3(24/2)
----
x = 3*12 = 36 (# of oranges you can buy for $5.25)
---------------------
Cheers,
Stan H.